OpenStudy (smiths21):

F(x) =10x+5 and g(x) =3x+6 then what is (f-g) (-1)??

OpenStudy (data_lg2):

Hi If you plug in x=-1 in the function f(x), what will you get? and if you plug in x=-1 in the function g(x), what will you get?

OpenStudy (smiths21):

thank you

OpenStudy (smiths21):

-5,3

OpenStudy (data_lg2):

Right.. Now, we have to subtract the function g(x) from f(x). (f-g) (-1) --> this is the same as : f(-1)-g(-1) f(-1)-g(-1) = -5 -3 = ...?